OpenStudy (anonymous):

3m - 5 = 7m + 3 m = ?

OpenStudy (ash2326):

\[3m-5=7m+3\] subtract both sides by 3m \[3m-3m-5=7m-3m+3\] \[-5=4m+3\] now subtract bots sides by 3 \[-5-3=4m+3-3\] \[-8=4m\] divide both sides by 4 \[-2=m\] so \[m=-2\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (preetha):

Good work Ash. Very nice explanation.

OpenStudy (ash2326):

Thanks preetha:)